16) Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of the necessary parts for its engines, including all of the carburetors. An outside supplier has offered to sell one type of carburetor to Troy Engines, Ltd., for a cost of $33 per unit. To evaluate this offer, Troy Engines, Ltd., has gathered the following information relating to its own cost of producing the carburetor internally:

Per Unit 18,000 Units
Per Year
Direct materials $ 15 $ 270,000
Direct labor 9 162,000
Variable manufacturing overhead 4 72,000
Fixed manufacturing overhead, traceable 6 * 108,000
Fixed manufacturing overhead, allocated 9 162,000
Total cost $ 43 $ 774,000

*One-third supervisory salaries; two-thirds depreciation of special equipment (no resale value).

Required:

1. Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 18,000 carburetors from the outside supplier?

2. Should the outside supplier’s offer be accepted?

3. Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $180,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 18,000 carburetors from the outside supplier?

4. Given the new assumption in requirement 3, should the outside supplier’s offer be accepted?

17) Imperial Jewelers manufactures and sells a gold bracelet for $403.00. The company’s accounting system says that the unit product cost for this bracelet is $261.00 as shown below:

Direct materials $ 141
Direct labor 87
Manufacturing overhead 33
Unit product cost $ 261

The members of a wedding party have approached Imperial Jewelers about buying 26 of these gold bracelets for the discounted price of $363.00 each. The members of the wedding party would like special filigree applied to the bracelets that would require Imperial Jewelers to buy a special tool for $453 and that would increase the direct materials cost per bracelet by $14. The special tool would have no other use once the special order is completed.

To analyze this special order opportunity, Imperial Jewelers has determined that most of its manufacturing overhead is fixed and unaffected by variations in how much jewelry is produced in any given period. However, $15.00 of the overhead is variable with respect to the number of bracelets produced. The company also believes that accepting this order would have no effect on its ability to produce and sell jewelry to other customers. Furthermore, the company could fulfill the wedding party’s order using its existing manufacturing capacity.

Required:

1. What is the financial advantage (disadvantage) of accepting the special order from the wedding party?

2. Should the company accept the special order?
Answer #1

16a) Differential analysis

Make Buy
Direct material 270000
Direct labor 162000
Variable manufacturing overhead 72000
Fixed manufacturing overhead, traceable (108000/3) 36000
Purchase cost (18000*33) 594000
Total relevant cost 540000 594000

Financial (disadvantage) = 540000-594000 = -54000

b) Rejected

c) Differential analysis

Make Buy
Direct material 270000
Direct labor 162000
Variable manufacturing overhead 72000
Fixed manufacturing overhead, traceable (108000/3) 36000
Opportunity Cost 180000
Purchase cost (18000*33) 594000
Total relevant cost 720000 594000

Financial advantage = 720000-594000 = 126000

d) Accepted

Note : Please post each question individually as per Chegg guidelines